What is 19-3x=28 im confused by the postition of the variable and how to solve it
Found 2 solutions by Alan3354, epena:
Answer by Alan3354(69443)

What is 19-3x=28
It's an equation.
---------------
im confused by the postition of the variable and how to solve it
19-3x=28
Change the position
-3x + 19 = 28
Subtract 19
-3x = 9
Divide by -3
x = -3

Answer by epena(3)

19−3x=28
Since 19 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 19 from both sides.
−3x=−19+28
Add 28 to −19 to get 9.
−3x=9
Divide each term in the equation by −3.
−3x −3 =9 −3
Move the minus sign from the denominator to the front of the expression.
3x 3 =9 −3
Cancel the common factor of 3 the expression 3x 3 .
1 3 x 3 =9 −3
Reduce the expression 3x 3 by removing a factor of 3 from the numerator and denominator.
x=9 −3
Move the minus sign from the denominator to the front of the expression.
x=−9 3
Cancel the common factor of 3 the expression −9 3 .
x=− 3 9 3
Reduce the expression −9 3 by removing a factor of 3 from the numerator and denominator.
x=−3